OpenStudy (anonymous):

Evaluate the integral: dx / e^(pie*x) from 2 (upper bound) to 0.

OpenStudy (anonymous):

\[\int\limits_{0}^{2}e ^{-pix}dx = -e ^{-2\pi}+1/\pi\]

OpenStudy (anonymous):

Why is it that? Also, what happened to the X?

OpenStudy (anonymous):

im looking at what Wolfram Alpha gave me, and their answer is (1/pi)e^(pi*x)

OpenStudy (anonymous):

this is an definite integral, we have to evaluate this primitive. You can make an substitution y = -pix and find this function

OpenStudy (anonymous):

ahhh im only confusing myself. I set u = -pi*x. du = -pi dx So, integral of e^(-pi*x) becomes e^(u) * dx. Taking the integral of that is just e^(u)

OpenStudy (anonymous):

yes. but excuse me I have a mistake evaluating function. Don't forget the limits of integration

OpenStudy (anonymous):

So it becomes: e^(-pi*x) from 2 to 0? and that's the answer? (plug in 2 for the equation, and plug in 0 and subtract the two)?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

okay thanks :)